Problem:

Vulcan Company's contribution format income statement for June is given below:

Vulcan Company
Income Statement
For the Month Ended June 30
  Sales $750,000  
  Variable expenses

336,000  

  Contribution margin 414,000  
  Fixed expenses

378,000  

  Net operating income

$36,000  


 

     Management is disappointed with the company's performance and is wondering what can be done to improve profits. By examining sales and cost records, you have determined the following:

 

a.

The company is divided into two sales territories—Northern and Southern. The Northern territory recorded $300,000 in sales and $156,000 in variable expenses during June; the remaining sales and variable expenses were recorded in the Southern territory. Fixed expenses of $120,000 and $108,000 are traceable to the Northern and Southern territories, respectively. The rest of the fixed expenses are common to the two territories.

b.

The company is the exclusive distributor for two products—Paks and Tibs. Sales of Paks and Tibs totaled $50,000 and $250,000, respectively, in the Northern territory during June. Variable expenses are 22% of the selling price for Paks and 58% for Tibs. Cost records show that $30,000 of the Northern territory's fixed expenses are traceable to Paks and $40,000 to Tibs, with the remainder common to the two products.


Question plus work done so far:

Prepare contribution format segmented income statements. (Round percentage computations to 1 decimal place. Omit the "$" and "%" signs in your response.)

 

      Sales Territory
  Total Company Northern Southern
  Amount % Amount % Amount %
  Sales $750,000  
100  
300,000$       $ 450,000
   
  Variable expenses

336,000

 

156,000

 

180,000

 

  Contribution margin

414,000

 

144,000

 

270,000

 

  Traceable fixed expenses

228,000

 

120,000

 

108,000

 

  Sales territory segment margin 186,000
   

$ 24,000

 

$ 162,000

 

  Common fixed expenses 150,000
           
  Net operating income

$36,000 

 

       

 
      Product Line
  Northern Territory Paks Tibs
  Amount % Amount % Amount %
  Sales $ 300,000
100  
$50,000 
    $250,000 
   
  Variable expenses

156,000

 

11,000

22

145,000

58

  Contribution margin 144,000
    39,000
    105,000
   
  Traceable fixed expenses

70,000

 

30,000

 

40,000

 

  Product line segment margin 74,000
   

$9,000 

 

$65,000 

 

  Common fixed expenses 50,000
           
  Sales territory segment margin

$ 24,000

 

       

 

 

My Question:

How do I get the percentage shown in each cell? the only percentage I got right so far are correct because they were given to me in the problem. I just dont know how to get the rest.

 

First person to explain how to get the % for the cells will receive life saver. Also the % cell for variable expenses for the second segment is not 80%.
Answer #1
Concepts and reason

Financial Statement: The statement is prepared for the specific period which comprises the financial information of the organization. It includes a statement of income which shows the profitability of the business, balance sheet which shows the financial position in the terms of assets, liabilities, and capital, and cash flow statement which represents the cash flows for the accounting period.

Fundamentals

Income Statement: It is a statement which is prepared periodical and shows the profitability of the business conducted for a particular period. Under this statement, all expenses, losses, incomes, and gains related to a particular period are recorded. Expenses and losses are debited in the income statement and on the other side, incomes and gains are credited to the Income Statement.

Segment analysis: Segmentation means to divide the market into parts, or segments, which are available, money-making and have a growth potential. In other words, it is quite possible for a company to target the entire market, because of cost, time and effort restrictions. Analyzing the various components may be done on the basis of percentage of total sales.

Sales: It refers the revenue which is earned by the core business of the organization. The units sold during the period are multiplied by the selling price per unit in order to determine the amount of sales.

Variable expenses: These are the expenses which can be directly traceable to the product are included in the variable expenses. The units sold during the period are multiplied by the variable cost per unit in order to determine the amount of variable expenses.

Contribution: The first profit of the company is a contribution. It is determined by subtracting the amount of variable expenses from the amount of sales. The contribution is the mandatory requirement of the organization, all indirect business expenses are dependent on this amount of contribution.

Traceable fixed expenses: The fixed expenses are those which do not change with the level of activity. If the company does not produce a single unit, then also the occurrence of these expenses is fixed. When fixed expenses can be classified between different lines of product, different departments on the specified basis, then those fixed expenses are called traceable fixed expenses.

Common fixed expenses: The fixed expenses which cannot be classified between various products manufactured, various departments individually are known as common fixed expenses. Such fixed expenses are adjusted from the total amount of profits and cannot be dealt individually.

Net operating income: It refers the profit earned by the organization during the particular period from the operating business of the organization. It is resulting in reducing the variable expenses and fixed expenses from the sales revenue generated during the period. Expenses are related to manufacturing and operations of the business.

(1a)

Prepare the Segmented Income Statement in form of contribution format:

Sales Territory
Total Company
Amount (S) Amount (S) Amount (S)
750,000
S
450,000
180,000
Sales
Variable expenses
Contribution

Working notes:

Total Company:

Compute the Contribution margin:

Contributionmargin=SalesVariableexpenses=$750,000$336,000=$414,000\begin{array}{c}\\{\rm{Contribution}}\,{\rm{margin}} = {\rm{Sales}} - {\rm{Variable}}\,{\rm{expenses}}\\\\ = \$ 750,000 - \$ 336,000\\\\ = \$ 414,000\\\end{array}

Compute the traceable fixed expenses:

Traceablefixedexpenses=TraceablefixedexpensesofSegmentN+TraceablefixedexpensesofSegmentS=$120,000+$108,000=$228,000\begin{array}{c}\\{\rm{Traceable}}\,\,{\rm{fixed}}\,{\rm{expenses}} = {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}}\,{\rm{of}}\,{\rm{Segment}}\,{\rm{N}}\\\\ + {\rm{Traceable}}\,\,{\rm{fixed}}\,{\rm{expenses}}\,{\rm{of}}\,{\rm{Segment}}\,{\rm{S}}\\\\ = \$ 120,000 + \$ 108,000\\\\ = \$ 228,000\\\end{array}

Compute Sales territory segment income:

Salesterritorysegmentincome=ContributionmarginTraceablefixedexpenses=$414,000$228,000=$186,000\begin{array}{c}\\{\rm{Sales}}\,{\rm{territory}}\,{\rm{segment}}\,{\rm{income}} = {\rm{Contribution}}\,{\rm{margin}} - {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 414,000 - \$ 228,000\\\\ = \$ 186,000\\\end{array}

Compute the common fixed expenses:

Commonfixedexpenses=TotalfixedexpensesTraceablefixedexpenses=$378,000$228,000=$150,000\begin{array}{c}\\{\rm{Common}}\,{\rm{fixed}}\,{\rm{expenses}} = {\rm{Total}}\,{\rm{fixed}}\,{\rm{expenses}} - {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 378,000 - \$ 228,000\\\\ = \$ 150,000\\\end{array}

Compute the net operating income:

NetOperatingIncome=SalesterritorysegmentincomeCommonfixedexpenses=$186,000$150,000=$36,000\begin{array}{c}\\{\rm{Net}}\,{\rm{Operating}}\,{\rm{Income}} = {\rm{Sales}}\,{\rm{territory}}\,{\rm{segment}}\,{\rm{income}} - {\rm{Common}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 186,000 - \$ 150,000\\\\ = \$ 36,000\\\end{array}

Sales Territory N:

Compute the Contribution margin:

Contributionmargin=SalesVariableexpenses=$300,000$156,000=$144,000\begin{array}{c}\\{\rm{Contribution}}\,{\rm{margin}} = {\rm{Sales}} - {\rm{Variable}}\,{\rm{expenses}}\\\\ = \$ 300,000 - \$ 156,000\\\\ = \$ 144,000\\\end{array}

Compute Sales territory segment income:

Salesterritorysegmentincome=ContributionmarginTraceablefixedexpenses=$144,000$120,000=$24,000\begin{array}{c}\\{\rm{Sales}}\,{\rm{territory}}\,{\rm{segment}}\,{\rm{income}} = {\rm{Contribution}}\,{\rm{margin}} - {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 144,000 - \$ 120,000\\\\ = \$ 24,000\\\end{array}

Sales Territory S:

Compute the Sales amount of Territory S:

SalesofTerritoryS=SalesoftotalcompanySalesofTerritoryN=$750,000$300,000=$450,000\begin{array}{c}\\{\rm{Sales}}\,\,{\rm{of}}\,{\rm{Territory}}\,{\rm{S}} = \,{\rm{Sales}}\,\,{\rm{of}}\,{\rm{total}}\,\,{\rm{company}} - {\rm{Sales}}\,\,{\rm{of}}\,{\rm{Territory}}\,{\rm{N}}\\\\ = \$ 750,000 - \$ 300,000\\\\ = \$ 450,000\\\end{array}

Compute the variable expenses of Territory S:

VariableexpensesofTerritoryS=VariableexpensesoftotalcompanyVariableexpensesofTerritoryN=$336,000$156,000=$180,000\begin{array}{c}\\{\rm{Variable}}\,\,{\rm{expenses}}\,\,{\rm{of}}\,{\rm{Territory}}\,{\rm{S}} = {\rm{Variable}}\,\,{\rm{expenses}}\,\,{\rm{of}}\,{\rm{total}}\,{\rm{company}}\\\\ - {\rm{Variable}}\,\,{\rm{expenses}}\,\,{\rm{of}}\,{\rm{Territory}}\,{\rm{N}}\\\\ = \$ 336,000 - \$ 156,000\\\\ = \$ 180,000\\\end{array}

Compute the Contribution margin:

Contributionmargin=SalesVariableexpenses=$450,000$180,000=$270,000\begin{array}{c}\\{\rm{Contribution}}\,{\rm{margin}} = {\rm{Sales}} - {\rm{Variable}}\,{\rm{expenses}}\\\\ = \$ 450,000 - \$ 180,000\\\\ = \$ 270,000\\\end{array}

Compute Sales territory segment income:

Salesterritorysegmentincome=ContributionmarginTraceablefixedexpenses=$270,000$108,000=$162,000\begin{array}{c}\\{\rm{Sales}}\,{\rm{territory}}\,{\rm{segment}}\,{\rm{income}} = {\rm{Contribution}}\,{\rm{margin}} - {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 270,000 - \$ 108,000\\\\ = \$ 162,000\\\end{array}

Analyze the Income Statement of the Company:

Total Company
Amount (S)%
750,000 100.0%
336,000 448%
414,000 55.2%
228,000 30.4%
186,000 24.8%
150,000 20.0%
36,000 48%
Sale

Working notes:

Analyze the Sales of the Company:

%ofSales=TotalSalesTotalSales=$750,000$750,000=100%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Total}}\,{\rm{Sales}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 750,000}}{{\$ 750,000}}\\\\ = 100\% \\\end{array}

Analyze the Variable expenses in relation to sales of the Company:

%ofSales=VariableexpensesTotalSales=$336,000$750,000=44.8%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Variable}}\,{\rm{expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 336,000}}{{\$ 750,000}}\\\\ = 44.8\% \\\end{array}

Analyze the Contribution margin in relation to sales of the company:

%ofSales=ContributionmarginTotalSales=$414,000$750,000=55.2%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Contribution}}\,{\rm{margin}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 414,000}}{{\$ 750,000}}\\\\ = 55.2\% \\\end{array}

Analyze the traceable fixed expenses in relation to sales of the company:

%ofSales=TraceablefixedexpensesTotalSales=$228,000$750,000=30.4%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Traceable fixed expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 228,000}}{{\$ 750,000}}\\\\ = 30.4\% \\\end{array}

Analyze the sales territory segment income in relation to sales of the company:

%ofSales=SalesterritorysegmentincomeTotalSales=$186,000$750,000=24.8%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Sales territory segment income}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 186,000}}{{\$ 750,000}}\\\\ = 24.8\% \\\end{array}

Analyze the common fixed expenses in relation to sales of the company:

%ofSales=CommonfixedexpensesTotalSales=$150,000$750,000=20%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Common fixed expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 150,000}}{{\$ 750,000}}\\\\ = 20\% \\\end{array}

Analyze the operating income in relation to sales of the company:

%ofSales=OperatingIncomeTotalSales=$36,000$750,000=4.8%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Operating}}\,{\rm{Income}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 36,000}}{{\$ 750,000}}\\\\ = 4.8\% \\\end{array}

Analyze the Income Statement of the Territory N:

Sales Territory
IN
Amount (S) %
300,000 100.0%
156,000 52.0%
144,000 48.0%
120,000 40.0%
Sales
Variable expenses
Contribution

Working notes:

Analyze the Sales of the Territory N:

%ofSales=TotalSalesTotalSales=$300,000$300,000=100%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Total}}\,{\rm{Sales}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 300,000}}{{\$ 300,000}}\\\\ = 100\% \\\end{array}

Analyze the Variable expenses in relation to sales of the Territory N:

%ofSales=VariableexpensesTotalSales=$156,000$300,000=52%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Variable}}\,{\rm{expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 156,000}}{{\$ 300,000}}\\\\ = 52\% \\\end{array}

Analyze the Contribution margin in relation to sales of the Territory N:

%ofSales=ContributionmarginTotalSales=$144,000$300,000=48%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Contribution}}\,{\rm{margin}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 144,000}}{{\$ 300,000}}\\\\ = 48\% \\\end{array}

Analyze the traceable fixed expenses in relation to sales of the Territory N:

%ofSales=TraceablefixedexpensesTotalSales=$120,000$300,000=40%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Traceable fixed expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 120,000}}{{\$ 300,000}}\\\\ = 40\% \\\end{array}

Analyze the sales territory segment income in relation to sales of the Territory N:

%ofSales=SalesterritorysegmentincomeTotalSales=$24,000$300,000=8%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Sales territory segment income}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 24,000}}{{\$ 300,000}}\\\\ = 8\% \\\end{array}

Analyze the Income Statement of the Territory S:

Sales Territory
S
|Amount (S)%
450,000 100.0%
180,000 40.0%
270,000 60.0%
108,000 24.0%
162,000 36.0%
Sales
Variable expenses

Working notes:

Analyze the Sales of the Territory S:

%ofSales=TotalSalesTotalSales=$450,000$450,000=100%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Total}}\,{\rm{Sales}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 450,000}}{{\$ 450,000}}\\\\ = 100\% \\\end{array}

Analyze the Variable expenses in relation to sales of the Territory S:

%ofSales=VariableexpensesTotalSales=$180,000$450,000=40%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Variable}}\,{\rm{expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 180,000}}{{\$ 450,000}}\\\\ = 40\% \\\end{array}

Analyze the Contribution margin in relation to sales of the Territory S:

%ofSales=ContributionmarginTotalSales=$270,000$450,000=60%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Contribution}}\,{\rm{margin}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 270,000}}{{\$ 450,000}}\\\\ = 60\% \\\end{array}

Analyze the traceable fixed expenses in relation to sales of the Territory S:

%ofSales=TraceablefixedexpensesTotalSales=$108,000$450,000=24%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Traceable fixed expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 108,000}}{{\$ 450,000}}\\\\ = 24\% \\\end{array}

Analyze the sales territory segment income in relation to sales of the Territory S:

%ofSales=SalesterritorysegmentincomeTotalSales=$162,000$450,000=36%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Sales territory segment income}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 162,000}}{{\$ 450,000}}\\\\ = 36\% \\\end{array}

Prepare and analyze the Income Statement:

Sales Territory
Total Company
N
Amount (S)
300,000 100.0%
156,000 52.0%
144,000 48.0%
120,000 40.0%
8.0%
Amount (S)
450,000 1

(1b)

Prepare the Segmented Income Statement of Territory N in form of contribution format:

Product Line
T
N Territory
Amount (S) Amount (S) Amount (S)
300,000
P
т
250,000
Sales
Variable expenses
Contribution margin
T

Working notes:

N Territory:

Compute the traceable fixed expenses:

Traceablefixedexpenses=TraceablefixedexpensesofProductP+TraceablefixedexpensesofProductT=$30,000+$40,000=$70,000\begin{array}{c}\\{\rm{Traceable}}\,\,{\rm{fixed}}\,{\rm{expenses}} = {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}}\,{\rm{of}}\,{\rm{Product}}\,{\rm{P}}\\\\ + {\rm{Traceable}}\,\,{\rm{fixed}}\,{\rm{expenses}}\,{\rm{of}}\,{\rm{Product}}\,{\rm{T}}\\\\ = \$ 30,000 + \$ 40,000\\\\ = \$ 70,000\\\end{array}

Compute Product line segment income:

Productlinesegmentincome=ContributionmarginTraceablefixedexpenses=$144,000$70,000=$74,000\begin{array}{c}\\{\rm{Product line }}\,{\rm{segment}}\,{\rm{income}} = {\rm{Contribution}}\,{\rm{margin}} - {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 144,000 - \$ 70,000\\\\ = \$ 74,000\\\end{array}

Compute the common fixed expenses:

Commonfixedexpenses=TotalfixedexpensesTraceablefixedexpenses=$120,000$70,000=$50,000\begin{array}{c}\\{\rm{Common}}\,{\rm{fixed}}\,{\rm{expenses}} = {\rm{Total}}\,{\rm{fixed}}\,{\rm{expenses}} - {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 120,000 - \$ 70,000\\\\ = \$ 50,000\\\end{array}

Product line P:

Compute the variable expenses:

Variableexpenses=Sellingprice×33%=$50,000×22%=$11,000\begin{array}{c}\\{\rm{Variable}}\,\,{\rm{expenses}} = {\rm{Selling}}\,{\rm{price}} \times 33\% \\\\ = \$ 50,000 \times 22\% \\\\ = \$ 11,000\\\end{array}

Compute the Contribution margin:

Contributionmargin=SalesVariableexpenses=$50,000$11,000=$39,000\begin{array}{c}\\{\rm{Contribution}}\,{\rm{margin}} = {\rm{Sales}} - {\rm{Variable}}\,{\rm{expenses}}\\\\ = \$ 50,000 - \$ 11,000\\\\ = \$ 39,000\\\end{array}

Compute Product line segment income:

Productlinesegmentincome=ContributionmarginTraceablefixedexpenses=$39,000$30,000=$9,000\begin{array}{c}\\{\rm{Product line }}\,{\rm{segment}}\,{\rm{income}} = {\rm{Contribution}}\,{\rm{margin}} - {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 39,000 - \$ 30,000\\\\ = \$ 9,000\\\end{array}

Product line T:

Compute the variable expenses:

Variableexpenses=Sellingprice×58%=$250,000×58%=$145,000\begin{array}{c}\\{\rm{Variable}}\,\,{\rm{expenses}} = {\rm{Selling}}\,{\rm{price}} \times 58\% \\\\ = \$ 250,000 \times 58\% \\\\ = \$ 145,000\\\end{array}

Compute the Contribution margin:

Contributionmargin=SalesVariableexpenses=$250,000$145,000=$105,000\begin{array}{c}\\{\rm{Contribution}}\,{\rm{margin}} = {\rm{Sales}} - {\rm{Variable}}\,{\rm{expenses}}\\\\ = \$ 250,000 - \$ 145,000\\\\ = \$ 105,000\\\end{array}

Compute Product line segment income:

Productlinesegmentincome=ContributionmarginTraceablefixedexpenses=$105,000$40,000=$65,000\begin{array}{c}\\{\rm{Product line }}\,{\rm{segment}}\,{\rm{income}} = {\rm{Contribution}}\,{\rm{margin}} - {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 105,000 - \$ 40,000\\\\ = \$ 65,000\\\end{array}

Analyze the Income Statement of the Territory N:

N Territory
|Amount (S)
300,000 100.0%
156,000 52.0%
144,000 48.0%
70,000 23.3%
74,000 24.7%
50,000 16.7%
%
Sales
Variable ex

Working notes:

Analyze the Sales of the Territory N:

%ofSales=TotalSalesTotalSales=$300,000$300,000=100%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Total}}\,{\rm{Sales}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 300,000}}{{\$ 300,000}}\\\\ = 100\% \\\end{array}

Analyze the Variable expenses in relation to sales of the Territory N:

%ofSales=VariableexpensesTotalSales=$156,000$300,000=52%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Variable}}\,{\rm{expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 156,000}}{{\$ 300,000}}\\\\ = 52\% \\\end{array}

Analyze the Contribution margin in relation to sales of the Territory N:

%ofSales=ContributionmarginTotalSales=$144,000$300,000=48%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Contribution}}\,{\rm{margin}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 144,000}}{{\$ 300,000}}\\\\ = 48\% \\\end{array}

Analyze the traceable fixed expenses in relation to sales of the Territory N:

%ofSales=TraceablefixedexpensesTotalSales=$70,000$300,000=23.3%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Traceable fixed expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 70,000}}{{\$ 300,000}}\\\\ = 23.3\% \\\end{array}

Analyze the product line segment income in relation to sales of the Territory N:

%ofSales=ProductlinesegmentincomeTotalSales=$74,000$300,000=24.7%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Product line segment income}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 74,000}}{{\$ 300,000}}\\\\ = 24.7\% \\\end{array}

Analyze the common fixed expenses in relation to sales of the Territory N:

%ofSales=CommonfixedexpensesTotalSales=$50,000$400,000=16.7%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Common fixed expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 50,000}}{{\$ 400,000}}\\\\ = 16.7\% \\\end{array}

Analyze the sales territory segment income in relation to sales of the Territory N:

%ofSales=SalesterritorysegmentincomeTotalSales=$24,000$300,000=8%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Sales territory segment income}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 24,000}}{{\$ 300,000}}\\\\ = 8\% \\\end{array}

Analyze the Income Statement of the Product line P:

Product Line
P
Amount (S)
50,000 100.0%
11,000 22.0%
39,000 78.0%
30,000 60.0%
9,000 18.0%
%
Sales
Variable expenses
Contribu

Working notes:

Analyze the Sales of the Product P:

%ofSales=TotalSalesTotalSales=$50,000$50,000=100%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Total}}\,{\rm{Sales}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 50,000}}{{\$ 50,000}}\\\\ = 100\% \\\end{array}

Analyze the Variable expenses in relation to sales of the Product P:

%ofSales=VariableexpensesTotalSales=$11,000$50,000=22%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Variable}}\,{\rm{expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 11,000}}{{\$ 50,000}}\\\\ = 22\% \\\end{array}

Analyze the Contribution margin in relation to sales of the Product P:

%ofSales=ContributionmarginTotalSales=$39,000$50,000=78%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Contribution}}\,{\rm{margin}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 39,000}}{{\$ 50,000}}\\\\ = 78\% \\\end{array}

Analyze the traceable fixed expenses in relation to sales of the Product P:

%ofSales=TraceablefixedexpensesTotalSales=$30,000$50,000=60%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Traceable fixed expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 30,000}}{{\$ 50,000}}\\\\ = 60\% \\\end{array}

Analyze the product line segment income in relation to sales of the Product P:

%ofSales=ProductlinesegmentincomeTotalSales=$9,000$50,000=18%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Product line segment income}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 9,000}}{{\$ 50,000}}\\\\ = 18\% \\\end{array}

Analyze the Income Statement of the Product T:

Product Line
T
|Amount (S) %
250,000 100.0%|
145,000 58.0%
105,000 42.0%
40,000 16.0%
65,000 26.0%
Sales
Variable expenses
Co

Working notes:

Analyze the Sales of the Product T:

%ofSales=TotalSalesTotalSales=$250,000$250,000=100%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Total}}\,{\rm{Sales}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 250,000}}{{\$ 250,000}}\\\\ = 100\% \\\end{array}

Analyze the Variable expenses in relation to sales of the Product T:

%ofSales=VariableexpensesTotalSales=$145,000$250,000=58%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Variable}}\,{\rm{expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 145,000}}{{\$ 250,000}}\\\\ = 58\% \\\end{array}

Analyze the Contribution margin in relation to sales of the Product T:

%ofSales=ContributionmarginTotalSales=$105,000$250,000=42%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Contribution}}\,{\rm{margin}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 105,000}}{{\$ 250,000}}\\\\ = 42\% \\\end{array}

Analyze the traceable fixed expenses in relation to sales of the Product T:

%ofSales=TraceablefixedexpensesTotalSales=$40,000$250,000=16%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Traceable fixed expenses}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 40,000}}{{\$ 250,000}}\\\\ = 16\% \\\end{array}

Analyze the product line segment income in relation to sales of the Product T:

%ofSales=ProductlinesegmentincomeTotalSales=$65,000$250,000=26%\begin{array}{c}\\{\rm{\% }}\,\,{\rm{of}}\,{\rm{Sales}} = \frac{{{\rm{Product line segment income}}}}{{{\rm{Total}}\,{\rm{Sales}}}}\\\\ = \frac{{\$ 65,000}}{{\$ 250,000}}\\\\ = 26\% \\\end{array}

Prepare and analyze the Income Statement:

Product Line
N Territory
Amount (S) %
300,000 100.0%
156,000 52.0%
144,000 48.0%
70,000 23.3%
74,000 24.7%
50,000 16.7%|
T
P

Ans: Part 1a

The net operating income of the total company broken down between sales territories is $36,000 representing 4.8% of the total company sales.

Part 1b

The segmented income for the N Territory broken down by product line is

$24,000 representing 8.0% of the N Territory sales.